# SignMusketeers: An Efficient Multi-Stream Approach for Sign Language Translation at Scale

### v1

@@ 3.2 Self-Supervised Sign Components pre-training
Our method has two training phases. The first stage is self-supervised, and aims to produce encoders that specialize in sign language facial expressions and hand gestures. Using the DINOv2-Small architecture [46, 53], we pre-train the two encoders separately. We initialize ViT-small student and teacher backbones (due to computational constraints, we could not use a ViT-Base backbone as in Rust et al. [9]) with the teacher weights from the original dinov2_vits14_reg checkpoint, while the linear heads are randomly initialized.
We largely follow the training protocols of the DINOv2 paper with the recommended 4 registers [53]. The input face/hand images are randomly transformed into 2 global and 8 local views, using a scale of 0.5 to 1.0, and 0.25 to 0.5, respectively. Then, all views (both global and local) are passed to the student network. The student features are normalized with a Softmax to obtain the score vectors $p_{s}$. The teacher network, on the other hand, only accepts global views as input. We use the Sinkhorn-Knopp centering algorithm [54] on the features from the teacher network to obtain $p_{t}$. We compute cross-entropy loss between $p_{s}$ and $p_{t}$ and use it to update the student network. The gradient backpropagation is disabled for the teacher network. The weights of the teacher network are updated with an exponential moving average (ema) of the student network weights. In addition, DINOv2 also masks out random patches of the input views to the student and computes a patch-level cross-entropy loss (iBOT loss [55]) between the masked student tokens and the corresponding visible teacher tokens. Fig. 2 illustrates this setup.
We do this self-supervised pre-training using 1 million face crops and hand crops (obtained as described in Section (§3.1)) independently. We use a base learning rate of $2\times 10^{-4}$ and batch size per GPU of 128 on 8 A6000Ada GPUs (i.e., an effective batch size of 1024). To account for our relatively small dataset, we follow the recommendation of Roth et al. [56] and adjust the number of iterations per pseudo-epoch and the number of pseudo-epochs, resulting in 5 effective epochs in total. In the KoLeoLoss we change the hyperparameter $\epsilon$ from $10^{-8}$ to $10^{-4}$ to avoid infinite loss values.
The resulting teacher networks are used as our face encoder (DINOv2-F) and hand encoder (DINOv2-H) for the next stage, which is supervised training.
@@ 3.3 Supervised Sign Language Translation
Given a video of $T$ frames with an associated written language translation, we first obtain channel crops as described in (§3.1). Each crop is passed to the relevant frozen encoder, which is obtained as described in section (§3.2). This results in 3 $T\times 384$ matrices for the face, left hand and right hand. These matrices are projected to $T\times 256$ via stream-specific linear layers. The normalized body posture vectors are transformed to a higher dimensionality, from $T\times 14$ to $T\times 128$ (via a linear layer trained from scratch). All four feature streams (face, left hand, right hand, and body posture) are then temporally concatenated and projected to $T\times 768$ (the input size of the T5 model) via another linear layer, also trained from scratch.
To summarize: In stage 1, we independently pre-train two (face and hand specific) image-to-vector feature extractors. In stage 2, we jointly train a human pose feature linear transformation layer and a single linear layer transformation for the concatenated four-stream features, and we fine-tune the T5 model for translation. Both stages are shown in fig. 1.
@@ 4 Experiments
As in other works [12, 9], we use the YouTube-ASL and How2Sign [10] datasets. YouTube-ASL contains roughly 600,000 clips, or roughly 700 hours, of ASL video with weakly aligned English text translations. [^4] Although Uthus et al. [12] report a total of 610,193 clips in YouTube-ASL, we were only able to retrieve 601,995 clips, presumably because some clips have been deleted between the time of the dataset’s creation and our retrieval. How2Sign consists of 31,128 / 1,741 / 2,322 clips for the training / validation / test sets.
For the self-supervised training of DINOv2-Hand and DINOv2-Face, due to computational constraints we limit ourselves to 1 million random face crops and 1 million random hand crops from YouTube-ASL. We note that the state-of-the-art method proposed by Rust et al. [9] sees about 50 millions frames during pre-training.
@@
Note that Uthus et al. [12] and Rust et al. [9] include an additional training schedule, YT $+$ H2S, which involves training on a mixture of YT and H2S.
For the supervised training stage, similarly to other works [12, 9], we initiliaze our T5 from a T5.1.1-Base pre-trained checkpoint. We use a batch size of 128 (16 per GPU running on 8 GPUs), with other hyper-parameters identical to Rust et al. [9]. Additionally, when further fine-tuning on How2Sign after training on YouTube-ASL, we perform an extra 5,000 steps of fine-tuning.
@@ 4.1 Comparison to prior work
Table 1 compares our models to prior results on the How2Sign ASL-English translation task, including two approaches that train on a combination of YouTube-ASL and How2Sign. First, we observe that our method consistently outperforms the approach of Uthus et al. [12] across various metrics. For example, we improve BLEU by 1.9 points when using the YT→H2S training schedule. This improvement provides evidence for the benefits of using learned features over pose estimator features/coordinates. When using the H2S-only supervised training schedule, our scores are 1.2 BLEU scores above the ones of Uthus et al. [12] on this same training schedule (1.2 BLEU vs. 2.4 BLEU).
@@
Comparing with Rust et al. [9], in the most restrictive H2S-only schedule, our method is far behind SSVP-SLT (by 9 BLEU points). However, in the best-performing schedule (YT→H2S), the gap between the two approaches reduces significantly, and our performance is just 0.4 BLEU points below that of Rust et al. [9]. [^5] It is important to note that the authors of  Rust et al. [9] report an even higher BLEU score of 15.5 for a method that trains an additional CLIP [57] model on English text, on a union of YouTube-ASL and How2Sign. We include this result in fig. 3 (top right). Such techniques might also improve our performance, but we are unable to do the experiment due to computation constraints. We note that the same trend holds for the method of Uthus et al. [12] when compared to SSVP-SLT. On H2S alone, there is a substantial gap between the two methods (+10.5 BLEU points), but on YT→H2S, the gap reduces to +2.4 BLEU points.
@@
Intrinsically, human pose estimation, as used in the method of Uthus et al. [12], is multi-stream in nature since it returns vectors (coordinates) describing different body parts, which are later concatenated. Thus, we suspect that multi-stream approaches, in general, do not perform well with small datasets in supervised training. We hypothesize that this might be because the features they return are entirely frame-level features and do not yet contain any information about how these frame-level features relate to each other. In contrast, SSVP-SLT pre-trained features already have some information about how frames relate to each other.
@@
Nevertheless, with a larger dataset during the supervised training stage, this lead of SSVP-SLT diminishes drastically compared to multi-stream approaches like Uthus et al. [12] and our SignMusketeers method. This suggests that as the amount of labeled training data increases, the advantage of pre-trained features that capture some temporal relationships becomes less pronounced, and multi-stream approaches can bridge the performance gap.
@@
We suspect that the 0.4 BLEU difference between our result and that of SSVP-SLT may be attributable not only to the model architecture, but also to other significant factors that vary between the two approaches. One key difference is the amount of data used during pre-training. Our method uses a sample of only 1.2% of the YouTube-ASL frames, while SSVP-SLT uses 50% of the YouTube-ASL frames (every video at a stride of 2). Additionally, due to computational constraints, we pre-train our model for only 5 epochs, whereas SSVP-SLT pre-trains for 800 epochs. As another comparison, at epoch 100, SSVP-SLT achieves a BLEU score of 12.5 (see table 1 SSVP-SLT ${}_{100}^{\text{YT$+$H2S}(50)}$ and fig. 3). In addition, we note that SSVP-SLT is pre-trained on both YouTube-ASL and How2Sign—the target domain dataset—while our model is pre-trained only on YouTube-ASL. Fig. 3 shows the performance-resources tradeoff for multiple models, showing that our approach (top left, in blue) surpasses SSVP-SLT by 1.7 BLEU points while training for 20 times fewer pre-training epochs and without utilizing the target domain dataset during pre-training.
[…]
@@
These observations suggest that our method, despite using significantly less pre-training data and fewer pre-training epochs, can achieve competitive performance compared to the state-of-the-art. Further investigation into the impact of pre-training data size and the number of pre-training epochs on the final performance could provide valuable insights into the efficiency and scalability of sign language translation models.
[…]
Method BLEU-1 BLEU-2 BLEU-3 BLEU GPU-Hrs PT % of Frames
Supervised training Schedule: H2S
Lin et al. [58] $14.9$ $7.3$ $3.9$ $2.2$
Tarrés et al. [59] $34.0$ $19.3$ $12.2$ $8.0$
Uthus et al. [12] $15.0$ $5.1$ $2.3$ $1.2$
[…]
Supervised training Schedule: YT
Uthus et al. [12] $20.9$ $10.4$ $6.1$ $4.0$
[…]
Supervised training Schedule: YT $+$ H2S
Uthus et al. [12] $36.3$ $23.0$ $16.1$ $11.9$
[…]
Supervised training Schedule: YT $\rightarrow$ H2S
Uthus et al. [12] $37.8$ $24.1$ $16.9$ $12.4$
SSVP-SLT ${}_{800}^{\text{YT}(50)}$ [9] $41.9$ $27.7$ $19.8$ $14.7$ $18768^{*}$
[…]
4.2 Ablations
@@
BLEU-1 BLEU-2 BLEU-3 BLEU
Uncrop $+$ original DINOv2 23.5 12.4 7.6 4.9
Crop $+$ original DINOv2 36.5 23.1 15.9 11.3
Crop $+$ pre-trained DINOv2 $+$ global frame 36.7 23.8 16.7 12.2
Crop $+$ pre-trained DINOv2 (Ours) 41.5 27.2 19.3 14.3
@@
5 Conclusion
In this paper, we propose SignMusketeers, a data- and compute-efficient method for sign language translation at scale. It uses a multi-stream encoding scheme that focuses on important parts of a signing video (facial expressions, hands, and body posture) and requires only individual frames during self-supervised pre-training, in contrast to prior work on pre-training that requires long duration videos. SignMusketeers achieves competitive performance with roughly 40×40$\times$40 × less data and 50×50$\times$50 × less computation than prior work. Our ablation studies provide evidence that learning separate feature encoders that specialize in different aspects of sign language (face, hand, and body) greatly boosts translation performance. Our results suggest that temporal information may not be necessary during pre-training. We hope that our affordable approach can make sign language translation accessible to a wider range of researchers and practitioners.

### v2

@@ 3.2 Self-Supervised Sign Components Pre-training
Our method has two training phases. The first stage is self-supervised, and aims to produce encoders that specialize in sign language facial expressions and hand gestures. Using the DINOv2-Small architecture Oquab et al. (2023); Darcet et al. (2024), we pre-train the two encoders separately. We initialize ViT-small student and teacher backbones (due to computational constraints, we could not use a ViT-Base backbone as in Rust et al. (2024)) with the teacher weights from the original dinov2_vits14_reg checkpoint, while the linear heads are randomly initialized.
We largely follow the training protocols of the DINOv2 paper with the recommended 4 registers Darcet et al. (2024). The input face/hand images are randomly transformed into 2 global and 8 local views, using a scale of 0.5 to 1.0, and 0.25 to 0.5, respectively. Then, all views (both global and local) are passed to the student network. The student features are normalized with a Softmax to obtain the score vectors $p_{s}$. The teacher network, on the other hand, only accepts global views as input. We use the Sinkhorn-Knopp centering algorithm Caron et al. (2020) on the features from the teacher network to obtain $p_{t}$. We compute cross-entropy loss between $p_{s}$ and $p_{t}$ and use it to update the student network. The gradient backpropagation is disabled for the teacher network. The weights of the teacher network are updated with an exponential moving average (ema) of the student network weights. In addition, DINOv2 also masks out random patches of the input views to the student and computes a patch-level cross-entropy loss (iBOT loss Zhou et al. (2022)) between the masked student tokens and the corresponding visible teacher tokens. Fig. 3 illustrates this setup.
We do this self-supervised pre-training using 1 million face crops and hand crops (obtained as described in Section (§3.1)) independently. We use a base learning rate of $2\times 10^{-4}$ and batch size per GPU of 128 on 8 A6000Ada GPUs (i.e., an effective batch size of 1024). To account for our relatively small dataset, we follow the recommendation of Roth et al. (2024) and adjust the number of iterations per pseudo-epoch and the number of pseudo-epochs, resulting in 5 effective epochs in total. In the KoLeoLoss we change the hyperparameter $\epsilon$ from $10^{-8}$ to $10^{-4}$ to avoid infinite loss values.
The resulting teacher networks are used as our face encoder (DINOv2-F) and hand encoder (DINOv2-H) for the next stage, which is supervised training. We note that for the face, we found that blurring the whole face directly hurts performance. However, a modified approach of first greying areas except the eyes and mouth, and then blurring the remaining image, performs similarly to training on the whole face crop. We therefore use this strategy, which aims to balance between keeping important facial attributes while also having some privacy-awareness.
@@ 3.3 Supervised Sign Language Translation
Given a video of $T$ frames with an associated written language translation, we first obtain channel crops as described in (§3.1). Each crop is passed to the relevant frozen encoder, which is obtained as described in (§3.2). This results in 3 $T\times 384$ matrices for the face, left hand and right hand. These matrices are projected to $T\times 256$ via stream-specific linear layers. The normalized body pose vectors are transformed to a higher dimensionality, from $T\times 14$ to $T\times 128$ (via a linear layer trained from scratch). All four feature streams (face, left hand, right hand, and body pose) are then temporally concatenated and projected to $T\times 768$ (the input size of the T5 model) via another linear layer, also trained from scratch.
To summarize: In stage 1, we independently pre-train two (face- and hand-specific) image-to-vector feature extractors. In stage 2, we jointly train a human pose feature linear transformation layer and a single linear layer transformation for the concatenated four-stream features, and we fine-tune the T5 model for translation. Both stages are shown in Figure 1.
@@ 4 Experiments
Method BLEU-1 BLEU-2 BLEU-3 BLEU GPU-Hrs PT % of Frames
Supervised training Schedule: H2S
Lin et al. (2023) $14.9$ $7.3$ $3.9$ $2.2$
Tarrés et al. (2023) $34.0$ $19.3$ $12.2$ $8.0$
Uthus et al. (2023) $15.0$ $5.1$ $2.3$ $1.2$
[…]
Supervised training Schedule: YT
Uthus et al. (2023) $20.9$ $10.4$ $6.1$ $4.0$
[…]
*Supervised training Schedule: YT $+$ H2S
Uthus et al. (2023) $36.3$ $23.0$ $16.1$ $11.9$
[…]
Supervised training Schedule: YT $\rightarrow$ H2S
Uthus et al. (2023) $37.8$ $24.1$ $16.9$ $12.4$
SSVP-SLT ${}_{800}^{\text{YT}(50)}$ Rust et al. (2024) $41.9$ $27.7$ $19.8$ $14.7$ $18768^{*}$
[…]
Table 1: Quantitative results on How2Sign. GPU-Hrs = GPU hours used during the entire training stage. PT % of frames = Percentage of YouTube-ASL frames used in the self-supervised pre-training stage. Adjusted by throughput ratio reported at https://lambdalabs.com/gpu-benchmarks. *We did not include YT + H2S experiments as this setting does not reflect the common paradigm of pre-training on a large dataset and fine-tuning on a different smaller dataset. The YT → H2S paradigm better represents adaptability to real-world use cases.
As in other work Uthus et al. (2023); Rust et al. (2024), we use the YouTube-ASL and How2Sign Duarte et al. (2021) datasets. YouTube-ASL contains roughly 600,000 clips, or roughly 700 hours, of ASL video with weakly aligned English text translations. [^4] Although Uthus et al. (2023) report a total of 610,193 clips in YouTube-ASL, we were only able to retrieve 601,995 clips, presumably because some clips have been deleted between the time of the dataset’s creation and our retrieval. How2Sign consists of 31,128 / 1,741 / 2,322 clips for the training / validation / test sets.
For the self-supervised training of DINOv2-Hand and DINOv2-Face, due to computational constraints we limit ourselves to 1 million random face crops and 1 million random hand crops from YouTube-ASL. We note that the state-of-the-art method proposed by Rust et al. (2024) sees about 50 million frames during pre-training.
@@
Note that Uthus et al. (2023) and Rust et al. (2024) include an additional training schedule, YT $+$ H2S, which involves training on a mixture of YT and H2S.
For the supervised training stage, similarly to other work Uthus et al. (2023); Rust et al. (2024), we initiliaze our T5 from a T5.1.1-Base pre-trained checkpoint. We use a batch size of 128 (16 per GPU running on 8 GPUs), with other hyper-parameters identical to Rust et al. (2024). Additionally, when further fine-tuning on How2Sign after training on YouTube-ASL, we perform an extra 5,000 steps of fine-tuning.
@@ 4.1 Comparison to prior work
Table 1 compares our models to prior results on the How2Sign ASL-English translation task, including two approaches that train on a combination of YouTube-ASL and How2Sign. First, we observe that our method consistently outperforms the approach of Uthus et al. (2023) across various metrics. For example, we improve BLEU by 1.9 points when using the YT→H2S training schedule. This improvement provides evidence for the benefits of using learned features over pose estimator features/coordinates. When using the H2S-only supervised training schedule, our scores are 1.2 BLEU scores above the ones of Uthus et al. (2023) on this same training schedule (1.2 BLEU vs. 2.4 BLEU).
@@
Comparing with the SSVP-SLT approach of Rust et al. (2024), in the most restrictive H2S-only schedule, our method is far behind (by 9 BLEU points). However, in the best-performing schedule (YT→H2S), the gap between the two approaches reduces significantly, and our performance is just 0.4 BLEU points below that of Rust et al. (2024). [^5] It is important to note that the authors of  Rust et al. (2024) report an even higher BLEU score of 15.5 for a method that trains an additional CLIP Radford et al. (2021) model on English text, on the union of YouTube-ASL and How2Sign. We include this result in Figure 2 (top right). Such techniques might also improve our performance, but we are unable to do the experiment due to computation constraints. We note that the same trend holds for the method of Uthus et al. (2023) when compared to SSVP-SLT. On H2S alone, there is a substantial gap between the two methods (+10.5 BLEU points), but on YT→H2S, the gap reduces to +2.4 BLEU points.
@@
Human pose estimation, as used in the method of Uthus et al. (2023), is inherently multi-stream since it returns vectors (coordinates) describing multiple body parts, which are later concatenated. We suspect that both of the multi-stream approaches do not perform well with small datasets in supervised training. We hypothesize that this might be because the features they return are entirely frame-level features and do not yet contain any information about how these frame-level features relate to each other. In contrast, SSVP-SLT pre-trained features already have some information about how frames relate to each other.
@@
Nevertheless, with a larger dataset during the supervised training stage, the lead of SSVP-SLT diminishes drastically compared to multi-stream approaches like Uthus et al. (2023) and our SignMusketeers method. This suggests that as the amount of labeled training data increases, the advantage of pre-trained features that capture some temporal relationships becomes less pronounced, and multi-stream approaches can bridge the performance gap.
@@
We suspect that the 0.4 BLEU difference between our result and that of SSVP-SLT may be attributable not only to the model architecture, but also to other significant factors that vary between the two approaches. One key difference is the amount of data used during pre-training. Our method uses a sample of only 1.2% of the YouTube-ASL frames, while SSVP-SLT uses 50% of the YouTube-ASL frames (every video at a stride of 2). Additionally, due to computational constraints, we pre-train our model for only 5 epochs, whereas SSVP-SLT pre-trains for 800 epochs. As another comparison, at epoch 100, SSVP-SLT achieves a BLEU score of 12.5 (see Table 1 SSVP-SLT ${}_{100}^{\text{YT$+$H2S}(50)}$ and Figure 2). In addition, we note that SSVP-SLT is pre-trained on both YouTube-ASL and How2Sign—the target domain dataset—while our model is pre-trained only on YouTube-ASL. Fig. 2 shows the performance-resources tradeoff for multiple models, showing that our approach (top left, in blue) surpasses SSVP-SLT by 1.7 BLEU points while training for 20 times fewer pre-training epochs and without utilizing the target domain dataset during pre-training.
@@
These observations suggest that our method, despite using significantly less pre-training data and fewer pre-training epochs, can achieve competitive performance compared to the state of the art. Further investigation into the impact of pre-training data size and the number of pre-training epochs on the final performance could provide valuable insights into the efficiency and scalability of sign language translation models.
We perform ablation studies to investigate three key components of the model design, as shown in Appendix A. First, pre-training DINOv2 specifically on hand and face crops proves beneficial, improving performance. Second, the multi-stream approach demonstrates significant advantages over using just the original frames, significantly boosting the scores. Finally, adding raw frames as a fifth stream alongside the existing features (perhaps suprisingly) degrades performance. See Appendix A for more details.
5 Conclusion
SignMusketeers is a data- and compute-efficient method for sign language translation at scale. It uses a multi-stream encoding scheme that focuses on important parts of a signing video (facial expressions, hands, and body pose) and requires only individual frames during self-supervised pre-training, in contrast to prior work on pre-training that requires long duration videos. SignMusketeers achieves competitive performance with roughly 40×40$\times$40 × less data and 50×50$\times$50 × less computation than prior work.
Limitations
This research has several key limitations. The most significant is that despite matching or exceeding previous studies’ results, the overall performance remains inadequate. Therefore, machine translation cannot yet reliably replace human sign language interpreters across diverse contexts. Another limitation stems from the much smaller training datasets compared to those typically used for developing self-supervised models in the speech and text domains. Additionally, since this study focuses solely on American Sign Language, we cannot determine whether our approach would work effectively for other sign languages, even though different sign languages share common features and communication methods. Also, in this work we focus only on the sign language translation task from video to text. It will be important to adapt this approach to the opposite direction (text to video) and other sign language processing tasks such as isolated sign language recognition, continuous sign language recognition, and fingerspelling detection. To overcome these challenges, future research should focus on building larger datasets and incorporating multiple sign languages into the training process.
Acknowledgment
We are grateful to Anand Bhattad, Ankita Pasad, Ju-Chieh Chou, and Chung-Ming Chien for their valuable suggestions throughout this project.
Appendix A Ablations
BLEU-1 BLEU-2 BLEU-3 BLEU
Uncrop $+$ orig. 23.5 12.4 7.6 4.9
Crop $+$ orig. 36.5 23.1 15.9 11.3
Crop $+$ pre-tr. 36.7 23.8 16.7 12.2
$+$ global frame
Crop $+$ pre-tr. 41.5 27.2 19.3 14.3
@@ Appendix A Ablations
Appendix B Additional Analysis and Discussion
B.1 Scaling Analysis
To better understand the scaling properties of our approach, we conduct additional experiments varying the amount of pre-training data. Table 3 shows how model performance changes as we increase the number of frames used during pre-training from 0.1M to 1.0M frames on both How2Sign and YouTube-ASL datasets.
Pre-training Frames H2S YT $\rightarrow$ H2S
0.1M 1.5 11.6
0.4M 1.5 -
0.8M 2.1 -
1.0M 2.4 14.3
Table 3: Analysis of model performance (BLEU scores) with varying amounts of pre-training data.
The results demonstrate consistent improvement in translation quality as we increase the amount of pre-training data, with gains of 0.9 and 2.7 BLEU points on How2Sign and YouTube-ASL respectively when scaling from 0.1M to 1.0M frames. This suggests that our method can effectively leverage additional pre-training data, though the gains may be bounded by the model capacity of the ViT-Small architecture used in our experiments.
B.2 Multi-Stream Contribution Analysis
To validate the importance of each information stream in our multi-stream architecture, we conduct ablation experiments using different combinations of streams. Table 4 presents the results on the How2Sign dataset.
Streams Used BLEU
Face only 0.6
Hands only 0.2
Upper body only 0.8
Hands + Upper Body 2.1
All streams (full model) 2.4
Table 4: Impact of different information streams on translation performance (BLEU scores) on How2Sign.
These results empirically demonstrate that each stream contributes meaningful information to the translation task. While the upper body stream provides the strongest individual signal (0.8 BLEU), the combination of streams yields substantially better performance (2.4 BLEU), supporting our hypothesis that different visual cues complement each other. The relatively low performance of individual streams also highlights the importance of our multi-stream approach in capturing the full range of linguistic information present in sign language videos.
B.3 Choice of Visual Backbone
We investigate different visual backbones for feature extraction. While prior work has established DINOv2’s superiority over ResNet architectures for sign language tasks, we conducted additional experiments comparing DINOv2 with MAE on the How2Sign dataset.
Backbone BLEU
MAE (whole frame pre-trained) 0.3
DINOv2 (whole frame pre-trained) 1.0
DINOv2 (original) 1.1
Table 5: Comparison of visual backbones on How2Sign using a whole-frame approach.
The results, shown in Table 5, demonstrate DINOv2’s effectiveness as a feature extractor for sign language translation. Notably, the performance difference between pre-trained and original DINOv2 is minimal when using whole frames, which motivated our exploration of the multi-stream approach.
B.4 Data Efficiency and Scaling
To understand the relationship between pre-training data volume and model performance, we evaluate our model at different stages of pre-training, as shown in Table 6.
Pre-training data Frames BLEU
10% 0.1M 1.5
100% 1.0M 2.4
Table 6: Impact of pre-training data volume on translation performance on How2Sign.
The improvement from 1.5 to 2.4 BLEU with increased pre-training data suggests that our approach can effectively utilize additional data. However, we note that these gains are achieved while using only 1.2% of the available YouTube-ASL frames for pre-training, highlighting the efficiency of our method.
Appendix C Hyperparameter Settings
The hyperparameter values used in our experiments are shown in Table 7 and Table 8.
parameter value
dino:
loss_weight 1.0
head_nlayers 3
head_hidden_dim 2048
koleo_loss_weight 0.1
ibot:
loss_weight 1.0
mask_sample_probability 0.5
head_bottleneck_dim 256
head_nlayers 3
head_hidden_dim 2048
train:
batch_size_per_gpu 128
saveckp_freq 20
seed 0
num_workers 10
OFFICIAL_EPOCH_LENGTH 100
cache_dataset true
centering sinkhorn_knopp
student:
arch vit_small
patch_size 14
drop_path_rate 0.4
layerscale 1.0e-05
drop_path_uniform false
pretrained_weights null
ffn_layer mlp
block_chunks 0
qkv_bias true
proj_bias true
ffn_bias true
num_register_tokens 4
interpolate_antialias false
interpolate_offset 0.1
teacher:
momentum_teacher 0.994
final_momentum_teacher 1
warmup_teacher_temp 0.04
teacher_temp 0.07
warmup_teacher_temp_epochs 30
epochs 100
weight_decay 0.04
weight_decay_end 0.2
base_lr 0.0002
lr 0.0
warmup_epochs 10
min_lr 1.0e-06
clip_grad 3.0
freeze_last_layer_epochs 1
scaling_rule sqrt_wrt_1024
patch_embed_lr_mult 0.2
layerwise_decay 0.9
adamw_beta1 0.9
adamw_beta2 0.999
crops:
global_crops_scale (0.5,1.0)
local_crops_number 8
local_crops_scale (0.25,0.5)
global_crops_size 224
local_crops_size 98
Table 7: SignMusketeers pre-training settings.
parameter value
model_name t5-1.1
target_language English
expand_time false
freeze_model false
freeze_adapter false
force_target_language false
batch_size_per_gpu 16
learning_rate 0.001
train_steps 5000
fp16 false
test_every 500
save_every 500
log_every 500
num_workers 4
optimizer adamw_torch
schedule cosine
weight_decay 1e-1
label_smoothing_factor 0.2
generation_max_length 128
generation_num_beams 5
grad_clipping 1.0
Table 8: Fine-tuning settings for How2Sign (H2S).
Appendix D MediaPipe Pose Landmark Indices
The MediaPipe pose landmark indices are defined in Figure 4.
Figure 4: MediaPipe pose landmark indices. Source: Google AI MediaPipe Documentation.